OpenStudy (anonymous):

(6x2 – 5x + 2) ÷ (3x + 2)

OpenStudy (zehanz):

\[\frac{ 6x^2-5x+2 }{ 3x+2 }\]What do you have to do with this?

OpenStudy (anonymous):

I'm not entirely sure.

OpenStudy (zehanz):

If you did a long division you would get:\[2x-3+\frac{ 8 }{ 3x+2 }\]

OpenStudy (anonymous):

Alright.. Thanks.

OpenStudy (zehanz):

Why do a long division? Well, you can see now that for very large x (pos or neg) this expression looks more and more like 2x-3, so y=2x-3 is a slant asymptote...

OpenStudy (anonymous):

Alright..